Axolotl:

A contractor is building a pool labeled ABCD on the plans. If AC = 15y − 7 and BD = 2y + 6, what value of y ensures the pool is a rectangle? −13 −1 1 13

Axolotl:

This is similar to the other question I posted Tagging: @jhonyy9

Axolotl:

Hey! ty for looking at my question jhonyy :)

Axolotl:

This is similar to the other question you just helped me with

Axolotl:

I think the answer is 1

jhonyy9:

 so what you know about diagonales of a rectangle ?

Axolotl:

AC = BD

jhonyy9:

perfect and now solve it for y

Axolotl:

ok

jhonyy9:

15y -7 = 2y +6 y = ?

Axolotl:

-Add seven to both sides -15y = 2y + 13 -Subtract 2y from both sides -Combine like terms -13y = 13 -Divide both sides by 13 Y = 1 Am I correct?

jhonyy9:

so than you wan check if 1 is correct just substitute it inside these expression of AC and BD and you need get the same results for every two cases

Axolotl:

@jhonyy9 wrote:
so than you wan check if 1 is correct just substitute it inside these expression of AC and BD and you need get the same results for every two cases
ok

jhonyy9:

AC = BD 15y -7 = 2y +6 ???

Axolotl:

15y − 7 = 2y + 6 15(1) − 7 = 2(1) + 6 15 − 7 = 2 + 6 8 = 8

jhonyy9:

congrats perfect do you understand it now the way logical ?

Axolotl:

@jhonyy9 wrote:
congrats perfect do you understand it now the way logical ?
yes. Yay! THANK YOU!!!

Axolotl:

I have one more question that I need help with and then I'm good :)

jhonyy9:

ok post it please

Axolotl:

sure thing